Derive an equation for interference of light for bright and dark fringes


1
Expert's answer
2021-12-07T09:51:38-0500

When superimposing two coherent waves


"E_1=E_{01}\cos (\omega t-\frac{2\pi}{\lambda }r_1+\alpha_1)"

"E_2=E_{02}\cos (\omega t-\frac{2\pi}{\lambda }r_2+\alpha_2)"


the intensity of the resulting wave is equal to "I" ~ "E^2" . So, we get

...

"I=I_1+I_2+2\sqrt{I_1I_2}\cos\Delta\phi" .


The equation shows that at points in space where "\cos\Delta\phi>0,\ I>I_1+I_2" , ie there is an increase in intensity (maximums). If "\cos\Delta\phi<0,\ I<I_1+I_2" . In this case there is a decrease in light intensity (minimums).


The wave intensity will be maximum at "\cos\Delta\phi=1" . Then:


"\Delta \phi=\pm2\pi k,\ \Delta=\pm k\lambda\ (k=0,1,2, ...)" .


The minimum intensity will correspond to "\cos\Delta\phi=-1" . In this case:


"\Delta \phi=\pm(2k+1)\pi,\ \Delta=\pm (2k+1)\frac{\lambda}{2}\ (k=0,1,2, ...)" ,


where "\Delta" is the optical path length.